The pendulum of a certain clock has time period 2.04 s. How fast or slow does the cock run during 24 hours?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. . The pendulum of the clock has time period 2.04sec. Now, No. or oscillation in 1 day = (24 × 3600)/2 = 43200 But, in each oscillation it is slower by (2.04 – 2.00) = 0.04sec. So, in one day it is slower by, = 43200 × (0.04) = 12 sec = 28.8 min So, the clock runs 28.8 minutes slower in one day.